OpenStudy (anonymous):

What is the area of a circle with a diameter of 22 meters? Use 22/7 for pi . choices: 374 m^2 370 6/7 m^2 380 2/7 m^2 377 1/7 m^2

OpenStudy (lgbasallote):

area = \(\large \frac{22}{7} (\frac{d}{2})^2\)

OpenStudy (lgbasallote):

so first divide your diameter by 2

Parth (parthkohli):

1) Get the value of the radius. 2) Use the formula: \(\Large \color{MidnightBlue}{\Rightarrow Area(Circle) = \pi r^2 }\)

OpenStudy (anonymous):

\[area=pi*r ^{2}\]